# Learning Stochastic Bridges for Video Object Removal via Video-to-Video Translation

### v1

@@ Abstract
Existing video object removal methods predominantly rely on diffusion models following a noise-to-data paradigm, where generation starts from uninformative Gaussian noise. This approach discards the rich structural and contextual priors present in the original input video. Consequently, such methods often lack sufficient guidance, leading to incomplete object erasure or the synthesis of implausible content that conflicts with the scene’s physical logic. In this paper, we reformulate video object removal as a video-to-video translation task via a stochastic bridge model. Unlike noise-initialized methods, our framework establishes a direct stochastic path from the source video (with objects) to the target video (objects removed). This bridge formulation effectively leverages the input video as a strong structural prior, guiding the model to perform precise removal while ensuring that the filled regions are logically consistent with the surrounding environment. To address the trade-off where strong bridge priors hinder the removal of large objects, we propose a novel adaptive mask modulation strategy. This mechanism dynamically modulates input embeddings based on mask characteristics, balancing background fidelity with generative flexibility. Extensive experiments demonstrate that our approach significantly outperforms existing methods in both visual quality and temporal consistency.
@@ 1 Introduction
In this work, inspired by recently proposed bridge models (Chen et al., 2021; Liu et al., 2023; Chen et al., 2023), we propose a paradigm shift by reformulating video object removal as a video-to-video translation task. We posit that the input video, even containing the unwanted object, holds a vast amount of valid structural and environmental information that should be leveraged as a strong prior rather than discarded in favor of Gaussian noise. To realize this, we introduce BridgeRemoval, a novel framework based on the stochastic bridge formulation (Tong et al., 2024; Zhou et al., 2023a; Chen et al., 2023) . Unlike standard diffusion-based methods that traverse a path from noise to data, our approach constructs a direct stochastic bridge between the source video distribution (the input video containing the object) and the target video distribution (the clean video with the object removed). By compressing videos into continuous latent representations via a Variational Autoencoder (VAE), we establish a tractable SDE-based generation process where the input latent representation serves directly as the prior boundary distribution.
@@
  • We propose BridgeRemoval, a novel video-to-video generative framework that utilizes an SDE-based bridge model to reformulate video object removal.
@@
  • Extensive experiments demonstrate the effectiveness of our approach in diverse real-world scenarios, outperforming existing methods in both visual quality and temporal coherence, while also highlighting the potential of our framework to be extended to other video-to-video edit tasks.
@@ 2 Related Works
Diffusion-based methods (Li et al., 2025b; Zi et al., 2025; Miao et al., 2025; Jiang et al., 2025; Lee et al., 2025) have achieved state-of-the-art performance in video object removal. DiffuEraser (Li et al., 2025b) adapts the image inpainting architecture of BrushNet (Ju et al., 2024) to the video domain via a progressive two-stage training strategy involving prior propagation and temporal refinement. Beyond simple occlusion handling, ROSE (Miao et al., 2025) addresses complex environmental interactions by leveraging a synthetic dataset with comprehensive supervision to simultaneously remove objects and their collateral visual effects, such as shadows and reflections. Furthermore, unified editing frameworks have also shown promise in this area. VACE (Jiang et al., 2025) proposes a universal backbone for diverse visual conditions, while GenOmnimatte (Lee et al., 2025) employs layered video decomposition, both of which support object removal as a downstream application.
@@ 3 Preliminaries
Probability path modeling (De Bortoli et al., 2021; Lipman et al., 2022; Liu et al., 2022) defines a class of generative models that describe continuous-time processes transporting mass from a prior distribution $p_{1}$ (source) to a target distribution $p_{0}$ (target). Generally, these models are represented by a stochastic differential equation (SDE) (Song et al., 2021):
@@ 3.1 Brownian Bridge
The standard Brownian Bridge (Albergo et al., 2023; Li et al., 2023) is a stochastic process conditioned to start at a specific point $x_{1}$ and end at $x_{0}$ within a fixed time interval. Unlike standard diffusion which maps data to Gaussian noise, a Brownian Bridge interpolates between two data points. Typically assuming a constant diffusion coefficient, the intermediate state $X_{t}$ at time $t$ can be written as a linear interpolation perturbed by noise with a fixed variance structure:
@@
Here, the variance term t(1t)t(1-t) is symmetric and reaches its maximum at $t=0.5$. While simple, this rigid variance schedule limits the model’s ability to handle complex transitions where the signal-to-noise ratio needs to be carefully controlled across time.
@@ 3.2 VP-SDE Bridge
To enable flexible control over the transport process, we adopt a Variance-Preserving (VP) formulation inspired by the VP-SDE typically used in score-based generative models (Song et al., 2021). We generalize the interpolation by introducing time-dependent coefficients derived from a noise schedule $\beta(t)\in[\beta_{\min},\beta_{\max}]$.
Given fixed endpoints $x_{0}$ (target) and $x_{1}$ (source), the intermediate state $X_{t}$ is defined by a weighted ternary interpolation:
@@
The coefficients $a_{t},b_{t},c_{t}$ are determined by the variance schedule $\sigma_{t}$:
@@
where $\sigma_{t}^{2}$ represents the cumulative variance at time $t$, $\sigma_{1}^{2}$ is the total variance at $t=1$, and $\bar{\sigma}_{t}^{2}=\sigma_{1}^{2}-\sigma_{t}^{2}$. This formulation guarantees boundary conditions satisfying $X_{0}=x_{0}$ and $X_{1}=x_{1}$. By adjusting $\beta(t)$, the VP-SDE Bridge provides a tunable trajectory that is well-suited for high-dimensional video generation tasks.
4 Methodology
@@ 4.2 Velocity-Driven Bridge Training
Standard diffusion models typically assume an uninformative Gaussian prior $\mathcal{N}(0,I)$. However, our task provides the source video $z_{\text{src}}$ as a strong structural prior. To leverage this, we adopt the Bridge Process formulation (Zhou et al., 2023a), replacing the Gaussian prior with a Dirac prior $\delta_{z_{\text{src}}}$.
Specifically, the forward process is modified from the standard diffusion SDE by introducing an additional drift term. The forward SDE is defined as:
@@
$+g(t)\mathrm{d}w,\quad z_{0}\sim p_{\text{data}},\ z_{T}=z_{\text{src}},$
where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $h(z_{t},t,z_{\text{src}})\triangleq\nabla_{z_{t}}\log p(z_{\text{src}}|z_{t})$ acts as a guidance term pulling the trajectory towards $z_{\text{src}}$.
Correspondingly, this process has a reverse process sharing the same marginal distribution, represented by the backward SDE:
@@
where $s_{\theta}(z_{t},t,z_{\text{src}})\triangleq\nabla_{z_{t}}\log p_{t,\text{bridge}}(z_{t}|z_{\text{src}})$ approximates the score function. Crucially, since the transition kernel remains Gaussian, the term $h$ is analytically tractable. The learning objective thus reduces to estimating the unknown score term $s_{\theta}$ via neural networks.
@@
Instead of simulating the stochastic paths directly via SDE solvers during training, we leverage the property that the marginal distribution of the VP-SDE Bridge at any time $t$ can be derived analytically. This allows us to train the model using a robust velocity-matching objective.
We sample a continuous timestep $t\in[0,1]$ and a noise variable $\epsilon\sim\mathcal{N}(0,I)$. The intermediate state $z_{t}$, following the marginal distribution of Eq. 7, is defined via a ternary interpolation:
$z_{t}=a_{t}z_{\text{tgt}}+b_{t}z_{\text{src}}+c_{t}\epsilon.$ (9)
Here, $z_{t}$ is a weighted mixture of the clean target $z_{\text{tgt}}$, the source prior $z_{\text{src}}$, and the noise $\epsilon$. The scalar coefficients $a_{t},b_{t},c_{t}$ are rigorously derived from the variance schedule $\sigma_{t}$ of the VP-SDE. Let $\sigma_{1}$ denote the total standard deviation at $t=1$, and $\bar{\sigma}_{t}=\sqrt{\sigma_{1}^{2}-\sigma_{t}^{2}}$. The coefficients are computed as:
@@
These coefficients ensure boundary conditions where $z_{0}=z_{\text{tgt}}$ and $z_{1}=z_{\text{src}}$.
To train the network $v_{\theta}$ parameterized by $\theta$, we define a ground-truth velocity target $u_{t}$, derived from the time-derivative of the bridge process (see Appendix):
@@ 4.3 Generative Inference via SDE Solver
During inference, we aim to reverse the bridge process to traverse from the source latent $z_{\text{src}}$ back to the clean target $z_{\text{tgt}}$. Theoretically, this corresponds to solving the backward SDE (Eq. 8) derived in Sec. 4.2.
Since the continuous backward SDE involves the unknown score function $s_{\theta}$, we employ a numerical SDE solver to approximate the trajectory. We adopt a variance-corrected sampling strategy that acts as a discretization of the backward SDE, ensuring the generation strictly follows the statistics of the Brownian bridge.
Given the current state $z_{t}$ at timestep $t$, the network predicts the velocity field $\hat{v}_{t}=v_{\theta}(z_{t},t,y)$, which approximates the drift term of the reverse process. We first estimate the clean target latent $\hat{z}_{0|t}$ by inverting the velocity equation (Eq. 11):
$\hat{z}_{0|t}=\frac{1}{\rho_{t}^{2}/c_{t}}\left(\frac{a_{t}}{c_{t}}z_{t}-\frac{a_{t}b_{t}}{c_{t}}z_{\text{src}}-\hat{v}_{t}\right).$ (13)
This formula explicitly removes the predicted noise and the weighted source component from $z_{t}$, recovering an approximation of the target $z_{\text{tgt}}$.
To progress to the next timestep $t^{\prime}<t$, we perform a solver step that interpolates between the current state $z_{t}$ and the predicted target $\hat{z}_{0|t}$, while injecting a calibrated amount of noise $\epsilon^{\prime}$:
$z_{t^{\prime}}=w_{1}z_{t}+w_{2}\hat{z}_{0|t}+w_{3}\epsilon^{\prime},\quad\text{with }\epsilon^{\prime}\sim\mathcal{N}(0,I).$ (14)
Here, $w_{1},w_{2},w_{3}$ are scalar weights analytically determined by the VP-SDE schedule to ensure the trajectory maintains the correct conditional variance. Specifically, $w_{3}$ modulates the contribution of the freshly sampled noise $\epsilon^{\prime}$ to prevent variance collapse near the target. The detailed inference algorithm is summarized in Algorithm 2.
@@ 4.4 Adaptive Mask Modulation
While the bridge formulation effectively anchors the generation trajectory to the source video, thereby ensuring superior background consistency, it introduces a critical trade-off. The source video acts as a strong prior, providing structural guidance that is beneficial for stability. However, this rigid adherence to the source distribution can become detrimental when removing large objects. In such cases, the strong structural prior from the source latent $z_{\text{src}}$ may overshadow the generative process, causing the model to struggle in deviating sufficiently from the original content to hallucinate plausible backgrounds for extensive occluded regions.
To mitigate this limitation without sacrificing the benefits of the bridge prior, we propose adaptive mask modulation. This mechanism is designed to dynamically recalibrate the feature influence based on the mask, enhancing the utilization of valid background context while relaxing the constraints on the inpainting region.
Let $h\in\mathbb{R}^{N\times D}$ denote the input patch embeddings to the Transformer, and $M\in\{0,1\}^{N\times 1}$ be the corresponding binary mask (where $1$ denotes the object region and 0 the background). We introduce a learnable modulation parameter $\gamma$, initialized to zero. The modulated embeddings $h^{\prime}$ are computed as:
@@
By adaptively boosting the magnitude of embeddings in the unmasked background regions, AMM effectively increases the “signal-to-noise” ratio of the valid context. In the self-attention mechanism, this amplification acts as a soft bias, encouraging the model to attend more heavily to the reliable background information when synthesizing the missing content. This strategy effectively counterbalances the strong source prior within the masked regions, providing the necessary generative flexibility to reconstruct large occlusions while maintaining strict fidelity to the surrounding environment.
@@ 5.1 Experimental Setup
Training Data. To ensure both geometric precision and semantic diversity, we construct a training dataset totaling approximately 47,000 video pairs from two complementary sources. First, we utilize the ROSE dataset (Miao et al., 2025), which consists of 16,000 synthetically generated samples. By leveraging 3D engines to toggle object visibility, ROSE provides perfectly aligned triplets of source videos, target backgrounds, and masks. While this offers high-quality supervision, our statistical analysis indicates a distribution bias towards static, rigid objects (e.g., signs, furniture), limiting the model’s generalization to complex motion. To address this limitation, we introduce a supplementary real-world composite dataset comprising 31,000 video pairs. This subset is explicitly designed to improve robustness on dynamic agents, containing 15,000 human-centric videos and 16,000 videos of common dynamic objects (e.g., animals, vehicles). The data pipeline adopts a composition-based approach: we employ SAM2 (Ravi et al., 2024) to extract high-fidelity foregrounds from real-world videos and composite them onto clean backgrounds. A Qwen3-VL (Bai et al., 2025) based filtering mechanism is then applied to ensure visual coherence and discard low-quality samples.

### v2

@@ Abstract
Existing video object removal methods predominantly rely on diffusion models following a noise-to-data paradigm, where generation starts from uninformative Gaussian noise. This approach discards the rich structural and contextual priors present in the original input video. Consequently, such methods often lack sufficient guidance, leading to incomplete object erasure or the synthesis of implausible content that conflicts with the scene’s physical logic. In this paper, we reformulate video object removal as a video-to-video translation task via a stochastic bridge model. Unlike noise-initialized methods, our framework establishes a direct stochastic path from the source video (with objects) to the target video (objects removed). This bridge formulation effectively leverages the input video as a strong structural prior, guiding the model to perform precise removal while ensuring that the filled regions are logically consistent with the surrounding environment. To address the trade-off where strong bridge priors hinder the removal of large objects, we propose a novel adaptive mask modulation strategy. This mechanism dynamically modulates input embeddings based on mask characteristics, balancing background fidelity with generative flexibility. Extensive experiments demonstrate that our approach significantly outperforms existing methods in both visual quality and temporal consistency. The project page is https://bridgeremoval.github.io/.
@@ 1 Introduction
In this work, inspired by recently proposed bridge models (Chen et al., 2021; Liu et al., 2023; Chen et al., 2023; Wang et al., 2025), we propose a paradigm shift by reformulating video object removal as a video-to-video translation task. We posit that the input video, even containing the unwanted object, holds a vast amount of valid structural and environmental information that should be leveraged as a strong prior rather than discarded in favor of Gaussian noise. To realize this, we introduce BridgeRemoval, a novel framework based on the stochastic bridge formulation (Tong et al., 2024; Zhou et al., 2023a; Chen et al., 2023) . Unlike standard diffusion-based methods that traverse a path from noise to data, our approach constructs a direct stochastic bridge between the source video distribution (the input video containing the object) and the target video distribution (the clean video with the object removed). By compressing videos into continuous latent representations via a Variational Autoencoder (VAE), we establish a tractable SDE-based generation process where the input latent representation serves directly as the prior boundary distribution.
@@
  • We propose BridgeRemoval, a novel video-to-video generative framework that utilizes an SDE-based bridge model to reformulate video object removal. By establishing a direct stochastic path from the source video to the target, our method effectively leverages input priors to achieve better object removal.
@@
  • To provide a comprehensive evaluation of video object removal models, we propose BridgeRemoval-Bench, which encompasses a wide variety of scenarios and provides meticulously annotated masks.
  • Extensive experiments demonstrate the effectiveness of our approach in diverse real-world scenarios, outperforming existing methods in both visual quality and temporal coherence, while also highlighting the potential of our framework to be extended to other video editing tasks.
@@ 2 Related Works
Diffusion-based methods (Li et al., 2025b; Zi et al., 2025a; Miao et al., 2025; Jiang et al., 2025; Lee et al., 2025) have achieved state-of-the-art performance in video object removal. DiffuEraser (Li et al., 2025b) adapts the image inpainting architecture of BrushNet (Ju et al., 2024) to the video domain via a progressive two-stage training strategy involving prior propagation and temporal refinement. Beyond simple occlusion handling, ROSE (Miao et al., 2025) addresses complex environmental interactions by leveraging a synthetic dataset with comprehensive supervision to simultaneously remove objects and their collateral visual effects, such as shadows and reflections. Furthermore, unified editing frameworks have also shown promise in this area. VACE (Jiang et al., 2025) proposes a universal backbone for diverse visual conditions, while GenOmnimatte (Lee et al., 2025) employs layered video decomposition, both of which support object removal as a downstream application.
@@ 3 Preliminaries
Probability path modeling (De Bortoli et al., 2021; Lipman et al., 2022; Liu et al., 2022) defines a class of generative models that describe continuous-time processes transporting mass from a prior distribution $p_{1}$ (source) to a target distribution $p_{0}$ (target). Generally, these models are governed by a stochastic differential equation (SDE) (Song et al., 2021):
@@ 3.1 Brownian Bridge
The standard Brownian Bridge (Albergo et al., 2023; Li et al., 2023) serves as a foundational instance of such processes, conditioned to start at a specific point $x_{1}$ and end at $x_{0}$. Unlike standard diffusion which maps data to Gaussian noise, a Brownian Bridge interpolates between two data points. Assuming a constant diffusion coefficient, the intermediate state $X_{t}$ follows a simple linear interpolation perturbed by noise:
@@
While intuitive, this symmetric variance schedule t(1t)t(1-t) is rigid. It lacks the flexibility to handle complex high-dimensional data like video, where the signal-to-noise ratio (SNR) needs to be carefully scheduled to ensure stable training and generation.
@@ 3.2 VP-SDE Bridge
To overcome the limitations of the standard Brownian Bridge, we adopt a Variance-Preserving (VP) formulation inspired by the VP-SDE typically used in score-based generative models (Song et al., 2021). This formulation generalizes the bridge process by introducing time-dependent coefficients derived from a flexible noise schedule $\beta(t)\in[\beta_{\min},\beta_{\max}]$.
Unlike the heuristic linear interpolation, the VP-SDE Bridge is theoretically grounded. As derived in Appendix B.1, the marginal distribution of this process remains Gaussian, allowing us to sample the intermediate state $X_{t}$ directly via a reparameterized ternary interpolation:
@@
Crucially, the coefficients $a_{t},b_{t},c_{t}$ are not arbitrary but are strictly determined by the cumulative variance schedule $\sigma_{t}$ of the VP-SDE:
@@
where $\sigma_{1}^{2}$ is the total variance at $t=1$, and $\bar{\sigma}_{t}^{2}=\sigma_{1}^{2}-\sigma_{t}^{2}$. This rigorous formulation guarantees that boundary conditions $X_{0}=x_{0}$ and $X_{1}=x_{1}$ are satisfied, while providing a tunable trajectory well-suited for high-fidelity video generation.
4 BridgeRemoval
@@ 4.2 Velocity-Driven Bridge Training
Standard diffusion models typically assume an uninformative Gaussian prior $\mathcal{N}(0,I)$. However, our task provides the source video $z_{\text{src}}$ as a strong structural prior. To leverage this, we adopt the Denoising Diffusion Bridge Model (DDBM) (Zhou et al., 2023a), essentially replacing the Gaussian prior with a Dirac prior $\delta_{z_{\text{src}}}$.
Specifically, the bridge process is defined by a forward SDE that drifts towards the source video $z_{\text{src}}$:
@@
$+g(t)\mathrm{d}w,\quad z_{0}\sim p_{\text{tgt}},\ z_{T}=z_{\text{src}},$
where $f(t)$ and $g(t)$ are the drift and diffusion coefficients of the VP-SDE. The term $h(z_{t},t,z_{\text{src}})\triangleq\nabla_{z_{t}}\log p(z_{\text{src}}|z_{t})$ acts as a guidance force, pulling the trajectory towards $z_{\text{src}}$.
Correspondingly, the generative reverse process is governed by:
@@
where $s_{\theta}(z_{t},t,z_{\text{src}})$ approximates the score function. Since the transition kernel remains Gaussian, the term $h$ is analytically tractable. We provide the proof of the Gaussian nature of the bridge marginal (see Appendix B.1) and the derivation of $h$ (see Appendix B.2).
@@
Instead of simulating stochastic paths directly, we leverage the analytic property of the bridge marginal distribution. As derived in Appendix B.1, the intermediate state $z_{t}$ can be sampled directly via a weighted ternary interpolation:
$z_{t}=a_{t}z_{\text{tgt}}+b_{t}z_{\text{src}}+c_{t}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$ (9)
Here, $z_{t}$ is a mixture of the clean target $z_{\text{tgt}}$, the source prior $z_{\text{src}}$, and noise $\epsilon$. The scalar coefficients $a_{t},b_{t},c_{t}$ are rigorously derived from the VP-SDE variance schedule $\sigma_{t}$ (see Appendix B.1):
@@
These coefficients ensure the boundary conditions $z_{0}=z_{\text{tgt}}$ and $z_{1}=z_{\text{src}}$.
To train the network $v_{\theta}$ parameterized by $\theta$, we define a ground-truth velocity target $u_{t}$. As shown in Appendix B.3, minimizing the velocity matching error is equivalent to score matching:
@@ 4.3 Generative Inference via SDE Solver
During inference, we aim to reverse the bridge process to traverse from the source latent $z_{\text{src}}$ back to the clean target $z_{\text{tgt}}$. Theoretically, this corresponds to solving the backward SDE (Eq. 8) derived in Sec. 4.2. Since the continuous backward SDE involves the unknown score function $s_{\theta}$, we employ a numerical SDE solver to approximate the trajectory. We adopt a variance-corrected sampling strategy that acts as a discretization of the backward SDE, ensuring the generation strictly follows the statistics of the Brownian bridge.
Given the current state $z_{t}$ at timestep $t$, the network predicts the velocity field $\hat{v}_{t}=v_{\theta}(z_{t},t,y)$, which approximates the drift term of the reverse process. We first estimate the clean target latent $\hat{z}_{0|t}$ by inverting the velocity equation (Eq. 11), utilizing the normalization factor $\rho_{t}=\sqrt{a_{t}^{2}+c_{t}^{2}}$ defined previously:
$\hat{z}_{0|t}=\frac{1}{\rho_{t}^{2}/c_{t}}\left(\frac{a_{t}}{c_{t}}z_{t}-\frac{a_{t}b_{t}}{c_{t}}z_{\text{src}}-\rho_{t}\hat{v}_{t}\right).$ (13)
This formula explicitly removes the predicted noise and the weighted source component from $z_{t}$, recovering an approximation of the target $z_{\text{tgt}}$. To progress to the next timestep $t^{\prime}<t$, we perform an SDE solver step that combines the current state $z_{t}$, the predicted target $\hat{z}_{0|t}$, and Gaussian noise injection:
$z_{t^{\prime}}=w_{1}z_{t}+w_{2}\hat{z}_{0|t}+w_{3}\epsilon^{\prime},\quad\epsilon^{\prime}\sim\mathcal{N}(0,I).$ (14)
Here, $w_{1},w_{2},w_{3}$ are scalar weights analytically determined by the bridge schedule to ensure the trajectory maintains the correct conditional variance. Specifically, $w_{3}$ modulates the contribution of the freshly sampled noise $\epsilon^{\prime}$ to prevent variance collapse near the target. The detailed inference algorithm is summarized in Algorithm 2.
@@ 4.4 Adaptive Mask Modulation
While the bridge formulation effectively anchors the generation trajectory to the source video, thereby ensuring superior background consistency, it introduces a critical trade-off. The source video acts as a strong prior, providing structural guidance that is beneficial for stability. However, this rigid adherence to the source distribution can become detrimental when removing large objects.
To mitigate this limitation, we propose adaptive mask modulation (AMM). This mechanism is designed to dynamically recalibrate the feature influence based on the mask, enhancing the utilization of valid background context while relaxing the constraints on the inpainting region.
Let $h\in\mathbb{R}^{N\times D}$ denote the input patch embeddings to the Transformer, and $M\in\{0,1\}^{N\times 1}$ be the corresponding binary mask (where $1$ denotes the object region and 0 the background). We introduce a learnable modulation parameter $\gamma$. The modulated embeddings $h^{\prime}$ are computed as:
@@
By adaptively boosting the magnitude of embeddings in the unmasked background regions, AMM effectively counterbalances the strong source prior within the masked regions, providing the necessary generative flexibility to reconstruct large occlusions while maintaining strict fidelity to the surrounding environment.
@@ 5.1 Experimental Setup
Training Data. We construct a training dataset totaling approximately 47,000 video pairs from two complementary sources. First, we utilize the ROSE dataset (Miao et al., 2025), which consists of 16,000 synthetically generated samples. By leveraging 3D engines to toggle object visibility, ROSE provides perfectly aligned triplets of source videos, target backgrounds, and masks. While this offers high-quality supervision, our statistical analysis indicates a distribution bias towards static, rigid objects, limiting the model’s generalization to complex motion. To address this limitation, we introduce a supplementary real-world composite dataset comprising 31,000 video pairs. This subset is explicitly designed to improve robustness on dynamic agents, containing 15,000 human-centric videos and 16,000 videos of common dynamic objects (e.g., animals, vehicles). The data pipeline adopts a composition-based approach: we employ SAM2 (Ravi et al., 2024) to extract high-fidelity foregrounds from real-world videos and composite them onto clean backgrounds. A Qwen3-VL (Bai et al., 2025b) based filtering mechanism is then applied to ensure visual coherence and discard low-quality samples.
Testing Data. Existing video object removal benchmarks suffer from limited scale and insufficient scene diversity. For instance, DAVIS (Perazzi et al., 2016) contains only 50 videos, all in landscape orientation, with some challenging cases involving fast motion. ROSE-Bench (Miao et al., 2025) includes 60 videos, but most depict static scenes, and its publicly released dataset does not specify a test split, making it unsuitable for standardized evaluation.
To enable comprehensive assessment of video object removal models, we introduce BridgeRemoval-Bench, a new benchmark comprising 150 high-quality videos with rich scene variety, including both landscape and portrait orientations, as well as annotations for both single and multiple objects. The dataset is constructed from three distinct sources:
  • 50 real-world videos collected from stock platforms such as Pexels (13), featuring aesthetically pleasing visuals and clean backgrounds.
  • 50 synthetic videos generated by first using Qwen3 (Yang et al., 2025a) to produce diverse text prompts and then synthesizing videos with Veo (12), covering common categories including humans, animals, and vehicles.
  • 50 in-the-wild videos sourced from social media platforms like TikTok (14), exhibiting highly complex and challenging scenarios.
For all 150 videos, object masks are generated using SAM2 (Ravi et al., 2024). Furthermore, to support methods such as VACE (Jiang et al., 2025) that require a textual description of the target video (object removed), we employ Qwen3-VL (Bai et al., 2025b) to generate a detailed caption for each video (object removed), and performed the same process on the DAVIS dataset. In summary, the DAVIS dataset and BridgeRemoval-Bench are used to evaluate the performance of various algorithms.
Metrics. To comprehensively evaluate the performance of each model, we employ the following metrics:
  • CLIP-T (Bai et al., 2025a) is used to calculate the similarity between the result video and the text description to verify whether the object has been removed. And CLIP-F (Bai et al., 2025a) calculates the average inter-frame CLIP similarity to gauge temporal consistency.
  • Following prior works (Bian et al., 2025), we evaluate the fidelity of the background using standard metrics, including PSNR, MSE, and LPIPS (Zhang et al., 2018). These metrics are computed exclusively on the unmasked areas.
  • We also report results from VBench (Huang et al., 2024). However, as these metrics from VBench exhibit limited discriminative power in our specific task and may not fully reflect the nuanced performance of the algorithms, we provide these results in the Appendix D for reference only.
Table 1: Quantitative comparisons among BridgeRemoval and other video object removal models in DAVIS dataset. Red stands for the best, Blue stands for the second best.
Video Quality Unmasked Region Preservation Method Year-Venue Base Model CLIP-T ↑ CLIP-F ↑ PSNR ↑ MSE ↓ LPIPS ↓ Runtime (s/frame) ↓ Senoria (Zi et al., 2025b) 2025-arxiv CogVideoX-5B (Yang et al., 2025b) 0.2618 0.9654 17.4752 2057.1134 0.4654 2.143 Ditto (Bai et al., 2025a) 2025-arxiv Wan2.1-14B (Wan et al., 2025) 0.2196 0.9712 11.7304 4945.1432 0.5226 6.402 ICVE (Liao et al., 2025) 2025-arxiv Hunyuanvideo-13B (Kong et al., 2025) 0.2361 0.9670 25.9675 244.8118 0.1116 12.571 ROSE (Miao et al., 2025) 2025-NeurIPS Wan2.1-1.3B (Wan et al., 2025) 0.2353 0.9676 26.1202 204.0729 0.1046 0.741 VACE (Jiang et al., 2025) 2025-ICCV Wan2.1-1.3B (Wan et al., 2025) 0.2629 0.9654 27.3758 138.2393 0.0796 1.765 GenOmnimatte (Lee et al., 2025) 2025-CVPR Wan2.1-1.3B (Wan et al., 2025) 0.2814 0.9701 26.6812 177.3141 0.125 3.914 BridgeRemoval (Ours) 2026 Wan2.1-1.3B (Wan et al., 2025) 0.2788 0.9768 28.2154 133.3605 0.0977 1.111
Table 2: Quantitative comparisons among BridgeRemoval and other video object removal models in BridgeRemoval-Bench. Red stands for the best, Blue stands for the second best.
Video Quality Unmasked Region Preservation Method Year-Venue Base Model CLIP-T ↑ CLIP-F ↑ PSNR ↑ MSE ↓ LPIPS ↓ Runtime (s/frame) ↓ Senoria (Zi et al., 2025b) 2025-arxiv CogVideoX-5B (Yang et al., 2025b) 0.2920 0.9843 18.9123 1553.9859 0.3365 2.143 Ditto (Bai et al., 2025a) 2025-arxiv Wan2.1-14B (Wan et al., 2025) 0.2218 0.9873 10.5340 7523.6043 0.5262 6.402 ICVE (Liao et al., 2025) 2025-arxiv Hunyuanvideo-13B (Kong et al., 2025) 0.2842 0.9887 17.4098 1589.3646 0.2740 12.571 ROSE (Miao et al., 2025) 2025-NeurIPS Wan2.1-1.3B (Wan et al., 2025) 0.2219 0.9879 30.5175 84.6950 0.0470 0.741 VACE (Jiang et al., 2025) 2025-ICCV Wan2.1-1.3B (Wan et al., 2025) 0.2875 0.9859 31.0548 61.7623 0.0335 1.765 GenOmnimatte (Lee et al., 2025) 2025-CVPR Wan2.1-1.3B (Wan et al., 2025) 0.2966 0.9878 31.5999 63.3032 0.0506 3.914 BridgeRemoval (Ours) 2026 Wan2.1-1.3B (Wan et al., 2025) 0.2922 0.9917 33.2394 46.3525 0.0415 1.111